OpenStudy (anonymous):

If you have a heat engine with a hot reservoir of 515°C and a cold reservoir of 190°C, what is the thermal efficiency of the system?

OpenStudy (anonymous):

someone help me

OpenStudy (michele_laino):

by definition the efficiency \( \large \eta \) of your termal machine is: \[\Large \eta = \frac{{{T_1} - {T_2}}}{{{T_1}}}\] where \( \large T_1, T_2 \) are the absolute temperatures., namely: \[\Large \begin{gathered} {T_1} \cong 515 + 273 = 788 \hfill \\ {T_2} \cong 190 + 273 = 463 \hfill \\ \end{gathered} \] so after a substitution, we get: \[\Large \eta = \frac{{{T_1} - {T_2}}}{{{T_1}}} = \frac{{788 - 463}}{{788}} = ...?\]